OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope and containing the midpoint of the segment whose end points are (2, -3) and (-6, 5) I tried 3 different times and none of them were right, help?

OpenStudy (anonymous):

I would start with y=mx+b. M being slope, b being y intercept. First find the slope of the two points \[(x_{1},y_{1}) (x _{2},y _{2})\] Then plug slope into y=mx+b. then plug in one of the points to solve for b. Understand?

OpenStudy (anonymous):

oops slope is 9/7

OpenStudy (anonymous):

and yeah i think so

OpenStudy (anonymous):

So if slope is 9/7. y=9/7x+b. You should plug in one of those points (x,y) and solve for b. once you get b plug it in and move the xs and ys to the same side for standard form

OpenStudy (anonymous):

actuallly, im not sure. the question itself confuses me. they use standard form, \[y-y _{1}=m\left( x - x_{1}\right) \]

OpenStudy (anonymous):

You just move x or y to one side of the equation once you get it in y=mx+b. y-mx=b

OpenStudy (anonymous):

\[(y-(-3)=m(x-2)\]with that just plug in one of points into the equation.